# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v3

@@ Abstract
We consider channels with synchronization errors modeled as insertions and deletions. A classical result for such channels is their information stability, hence the existence of the Shannon capacity, when the synchronization errors are memoryless. In this paper, we extend this result to the case where the insertions and deletions have memory. Specifically, we assume that the synchronization errors are governed by a stationary and ergodic finite state Markov chain, and prove that such channel is information-stable, which implies the existence of a coding scheme which achieves the limit of mutual information. This result implies the existence of the Shannon capacity for a wide range of channels with synchronization errors, with different applications including DNA storage. The methods developed may also be useful to prove other coding theorems for non-trivial channel sequences.
@@ I Introduction
A mathematical model of a physical channel is defined by conditional distributions $W(y|x)$ for all $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, which means that if a symbol $x\in{\mathcal{X}}$ is input, output $y\in{\mathcal{Y}}$ is observed with probability [^2] This is valid only for a countable ${\mathcal{Y}}$; for uncountable case, there should be output PDFs instead; in this paper uncountable ${\mathcal{Y}}$ are not considered. $W(y|x)$. The fact that $W$ is a channel with the input set ${\mathcal{X}}$ and the output set ${\mathcal{Y}}$ will be denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$. Shannon’s theorem [1] considers a memoryless channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ that is used $n$ times independently. The equivalent channel that transmits $n$ symbols simultaneously is the channel $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$, where the conditional probabilities are given by $W^{n}(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, with the use of the vector notation $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$.
In this paper, our focus is on channels with synchronization errors, modeled as insertions, deletions and channel errors. As an example, consider an insertion channel $V_{n}$ with $n$ input symbols $x\in{\mathcal{X}}^{n}$. In general, for each input symbol $x_{i}\in{\mathcal{X}}$, an output vector $y_{i}$ over ${\mathcal{X}}$ of length $\geq 1$ is produced. In contrast with the channel $W^{n}$, the resulting output of $V_{n}$ is a concatenation of output vectors, but not the vector $y_{1}^{n}$ of output vectors. For example, if input bits are $(1,0,0)$ and the corresponding outputs for each bit are $(1,0)$, $(0)$ and $(1,0)$, then the overall channel output would be $(1,0,0,1,0)$, not $((1,0),(0),(1,0))$.
In general, a communication setup, besides conditional probabilities of the original channel, also includes some underlying method of combining the outputs of multiple channel uses. In the case of simple memoryless channels, the combining method is just stacking together output symbols into the output vector; in the case of insertion (and/or deletion) channel, the combining method is concatenation. For this reason, in the non-trivial setup a general case of a channel sequence $(W_{n}:{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{n})_{n\in{\mathbb{N}}}$ is usually considered [2, 3], where sets ${\mathcal{X}}_{n}$, ${\mathcal{Y}}_{n}$ can vary for different $n$. For simplicity, we consider the case when input set for $W_{n}$ is ${\mathcal{X}}_{n}={\mathcal{X}}^{n}$ and ${\mathcal{X}}$ is fixed for all $n$.
For a channel sequence the mutual information capacity can be defined as $\lim_{n\to\infty}{\mathbf{I}}(W_{n})/n$, where ${\mathbf{I}}(W_{n})$ denotes the maximum mutual information between the input and the output of a channel $W_{n}$ over all possible input distributions. The coding capacity (or operational capacity) is the (maximum) asymptotic rate of coding schemes which achieve arbitrary small error probability as $n\to\infty$. The capacity theorem is a statement that the mutual information capacity is equal to the coding capacity.
There are many capacity theorems that generalize the Shannon’s theorem on non-trivial channel sequences. Dobrushin’s work [4] considers the most general case, where input and output alphabets can be continuous, and the metric of closeness of decoder’s output to the original signal can be arbitrary. It is shown that the so-called “information stability” of a channel sequence is sufficient for the capacity theorem to hold. Later, in [2] it was shown that information stability is necessary and sufficient for the capacity theorem. In [5], Dobrushin proves the capacity theorem for channels with independent synchronization errors. In that model, for each input symbol, an output vector of finite expected length is produced, and all the output vectors are concatenated to form the resulting output. Since this channel model includes any insertion/deletion/substitution errors, we call such model an IDS channel. The generalization of the capacity theorem to the case of IDS channels with continuous input alphabet is studied in [6]. Capacity theorems are also proven for finite-state Markov channels (FSMC) [7], for Gilbert-Elliott channels [8], for deletion channels concatenated with FSMC [9], for IDS channels with intersymbol interference [10].
IDS channels are used to model various communication channels, including bit patterned media recording [11] and DNA storage channels [12]. For the case of memoryless insertion/deletion channels, capacity upper bounds and lower bounds are also obtained in [13, 14, 15], and the asymptotic expansion of the independent identically distributed deletion channel capacity is found in [16]. See [17, 18] for a detailed review of progress in this area. While most existing results focus on the scenario with independent synchronization errors, it is also important to study the case with insertions/deletions with memory. For instance, current reading and writing technologies in DNA sequencing introduce memory to the channel errors [19, 20].
@@
In this paper, we consider IDS channels with Markov memory, i.e., the underlying insertions/deletions are governed by a Markov chain. This model is referred to as a Markov-IDS channel. That is, each state of the Markov chain is an IDS channel, called component channel, and the overall output of Markov-IDS channel is obtained by concatenating the individual component channel outputs. The Markov-IDS channel is a generalization of both IDS and FSMC: it is more general than IDS since statistics of the synchronization errors change over time according to Markov chain rule; it is more general then FSMC since it concatenates the outputs of each component IDS channel (which can be translated to non-concatenated version via inserting a special symbol at the end of each component channel’s output). We note that while this model generalizes the synchronization errors as studied by Dobrushin, it is not sufficiently general to encompass some DNA channel models with insertions/deletions with memory as in [19, 20] since the channel states do not depend on the channel input in our model. Alternatively, in [10], the capacity theorem is proven for IDS channels with intersymbol interference (ISI). The difference between the ISI-IDS channel model in [10] and Markov-IDS channel model in this paper is that the state of ISI-IDS channel is defined by a few recent input bits, and the channel state in the Markov-IDS model is random (though the distribution is defined by the previous state).
In this paper, we also introduce a unified notation, which helps to elucidate the idea behind Dobrushin’s methods, which can be seen as applying functions to channel sequences. Namely, if a function is applied on a channel’s output, and this function is not “doing much” in a certain sense, then the modified channel has the same mutual information capacity, coding capacity and information stability, as the original channel. This means that one can apply any finite number of such functions and still have the same capacities.
The paper is organized as follows. In Section II, notations and information-theoretic definitions are introduced. In Section III, functions on channels and their properties are described. In Section IV, the Markov-IDS channel sequence is defined. In Section V, the capacity theorem for Markov-IDS channel sequence is proven. In Section VI, conclusions are drawn.
II Background
II-A Notations
@@
We use the following notation. Sets are written in calligraphic letters ${\mathcal{X}},{\mathcal{Y}},{\overline{\mathcal{Y}}}$; operators are written in bold letters ${\mathbf{C}},{\mathbf{E}},{\mathbf{I}}$. For $b,c\in{\mathbb{N}}$, we denote a vector $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$. Here, elements $a_{i}$ can be of any type: a number, a distribution, a channel, etc. Also, we denote infinite sequences as $a^{\infty}=(a_{1},a_{2},\dots)$. For a set ${\mathcal{A}}$, we denote a set of all finite sequences with elements from ${\mathcal{A}}$ by ${\overline{\mathcal{A}}}={\left\{{()}\right\}}\cup\bigcup_{i=1}^{\infty}{\mathcal{A}}^{i}.$ For a vector $a=a_{1}^{n}\in{\overline{\mathcal{A}}}$ we denote by ${\mathbf{N}}(a)=n$ the vector length. We denote the concatenation (gluing) operator $\mathbf{g}$ as $\mathbf{g}(a_{1}^{n},b_{1}^{m})=(a_{1},a_{2},...,a_{n},b_{1},...,b_{m}),$ which is also generalized onto any number of input vectors. Also, define the function $\mathbf{g}_{n}:{\overline{\mathcal{A}}}^{n}\to{\overline{\mathcal{A}}}$, which glues $n$ vectors ${\overline{a}}_{i}\in{\overline{\mathcal{A}}}$ as $\mathbf{g}_{n}({\overline{a}}_{1}^{n})=\mathbf{g}({\overline{a}}_{1},{\overline{a}}_{2},...,{\overline{a}}_{n}).$ The inverse image $\mathbf{g}^{-1}_{n}(a)$ is the set of all possible ways of cutting $a\in{\overline{\mathcal{A}}}$ into $n$ subvectors, i.e., $\mathbf{g}^{-1}_{n}(a)={\left\{{{\overline{a}}_{1}^{n}|\mathbf{g}_{n}({\overline{a}}_{1}^{n})=a}\right\}}$. For a set ${\mathcal{S}}$, we denote by ${\mathbb{D}}_{\mathcal{S}}$ the set of all distributions over ${\mathcal{S}}$, i.e.
${\mathbb{D}}_{\mathcal{S}}={\left\{{f:{\mathcal{S}}\to[0,1]\left|\sum_{x\in{\mathcal{S}}}f(x)=1\right.}\right\}}.$
For $n\in{\mathbb{N}}$, denote by ${\mathbb{D}}_{n}$ the set of distributions over the set ${\left\{{1,2,\dots,n}\right\}}$ in a form of row vectors $\rho_{1}^{n}$, $\sum_{i=1}^{n}\rho_{i}=1$, $\rho_{i}\geq 0$.
In general, probabilities are usually defined using sigma-algebras. In this paper, however, all probabilities are defined on the countable sets, so the probability measure is always completely defined by the probabilities of the elementary events $\Pr{\left[{X=x}\right]}$.
@@
The probability distribution for any probability, expectation or entropy is written as a subscript. The values that are not known are interpreted as random variables, and it is assumed that we take summation/expectation over all their possible values. For example, if we have a conditional distribution $W(u,x|v,y)$, where $u\in{\mathcal{U}},v\in{\mathcal{V}},x\in{\mathcal{X}},y\in{\mathcal{Y}}$, and the values of $u$ and $v$ have been defined previously, then
$\Pr_{W(u,x|v,y)}{\left[{x>0}\right]}$ $=\sum_{\begin{subarray}{c}x\in{\mathcal{X}}:x>0\\ y\in{\mathcal{Y}}\end{subarray}}W(u,x|v,y),$
${\mathbf{E}}_{W(u,x|v,y)}{\left[{f(x,y)}\right]}$ $=\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(u,x|v,y)\cdot f(x,y).$
@@
${\mathbf{H}}_{W(x,y)}{\left[{W(x|y)}\right]}$ $=-\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(x,y)\cdot\log_{2}W(x|y),$
${\mathbf{H}}{\left[{W(x,y)}\right]}$ $=-\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(x,y)\cdot\log_{2}W(x,y).$

### v5

# Markov Insertion/Deletion Channels: Information Stability and Capacity Bounds

@@ Abstract
Particularly motivated by DNA storage channels, we consider channels with synchronization errors modeled as insertions and deletions, along with substitutions. We focus on the case where the synchronization error process has memory and investigate the information stability of these channels, hence the existence of their Shannon capacity. We assume that the synchronization errors are governed by a stationary and ergodic finite state Markov chain and prove that such a channel is information-stable, which implies the existence of a coding scheme that achieves the limit of mutual information. This result implies the existence of the Shannon capacity for a wide range of channels with synchronization errors, with different applications, including DNA storage. We also provide specific examples of deletion channels with Markov memory and numerically evaluate their capacity bounds, thereby allowing us to quantify the capacity difference between memoryless deletion channels and those with memory with the same deletion probability and reveal that having memory increases the channel capacity.
@@ I Introduction
Channels with synchronization errors are observed in different communication systems. For instance, for a wireless communication system at low signal-to-noise ratios, loss of synchronization results in cycle slips, causing mistimed sampling of the matched filter output, resulting in possible deletions, insertions, and substitutions. Hence, we refer to channels with synchronization errors as insertion/deletion/substitution (IDS) channels. As another example, misalignment of the read-head and write-head may cause deletion or insertion errors for bit-patterned media recording systems [1]. The insertion/deletion errors are in addition to other possible errors due to noise, modeled as an additive white Gaussian noise process or possibly asymmetric substitution errors. The emerging area of DNA storage represents another scenario in which IDS channels are observed in a practical setting [2].
There is extensive literature studying different aspects of transmission over channels with insertions/deletions. For instance, there are a number of number theoretic and algebraic code designs (suited for relatively short transmissions) (see, for example, [3, 4, 5] for some recent examples) and applications of concatenated coding solutions with inner marker or watermark codes [6, 7, 8]. On a different front, these channels have also been studied from an information-theoretic point of view. In particular, Dobrushin [9] proved that if the synchronization errors are memoryless – meaning that while for transmission of each channel input, the number of output symbols is not the same (hence the receiver cannot immediately identify which output symbols correspond to which input symbol), these errors are independent and identically distributed for all channel uses – the synchronization error channel is information stable, and hence the channel capacity exists. Specifically, the capacity expression is given by $C=\mathop{\lim}_{n\rightarrow\infty}\mathop{\max}_{p(\bf{X})}\frac{I(\bf{X};\bf{Y})}{n}$ where $\bf{X}$ is the length-$n$ channel input, and $\bf{Y}$ is the channel output, $I(\cdot;\cdot)$ is the mutual information between the input and the output sequences, and the maximization is over all stationary and ergodic processes.
While it is known that the Shannon capacity exists for memoryless synchronization error channels, it has been formidable to evaluate the above expression and compute the capacity, even for some seemingly simple models. For instance, the capacity of the independent and identically distributed (i.i.d.) deletion channel is still unknown. There are, however, upper and lower bounds on the capacity of different synchronization error channels developed over the years. For instance, Refs. [10, 11, 12] develop lower bounds on the capacity of i.i.d. deletion channels and channels with i.i.d. deletions and duplications. On the other hand, different works including [13, 14, 15] develop upper bounds on the capacity of i.i.d. deletion channels as well as channels with deletions/insertions/substitutions. Asymptotic expansion of the i.i.d. deletion channel capacity is determined in [16, 17]. Excellent surveys of the information-theoretic developments on channels with synchronization errors can be found in [18, 19].
Based on experimental results, it has recently been demonstrated that the insertion/deletion processes in DNA storage channels exhibit memory [20, 21]. In other words, the assumption of i.i.d. insertions and/or deletions falls short of some realistic channel models. Motivated by this fact, in this paper, we consider IDS channels with memory; specifically, we study the case where the synchronization errors are governed by an underlying Markov chain, referred to as Markov-IDS channels. That is, each state of the Markov chain is an IDS channel, called a component channel, and the overall output of the Markov-IDS channel is obtained by concatenating the individual component channel outputs. As the insertion/deletion processes have memory, Dobrushin’s result is not directly applicable. Hence, we need a new coding theorem to prove that the Shannon capacity of such channels exists, which is the main problem addressed in the paper. To prove our results, we extract different steps in the proof of the capacity theorem for synchronization error channels as independent propositions to allow for their use in the future in a streamlined manner. Furthermore, we provide several new results, stated as general propositions, on the perturbation of an optimal input distribution for a channel and the order of growth of the variance of mutual information density in connection to the smallest non-zero channel transition probability.
@@
Note that the Markov-IDS channel is a generalization of both IDS and Finite State Markov Channels (FSMC): it is more general than IDS since statistics of the synchronization errors change over time according to Markov chain rule; it is more general then FSMC since it concatenates the outputs of each component IDS channel (which can be translated to non-concatenated version via inserting a special symbol at the end of each component channel’s output). We highlight that the most closely related work to our paper is [22], which proves the capacity theorem for IDS channels with intersymbol interference (ISI). The difference between the ISI-IDS channel model in [22] and the Markov-IDS channel model in this paper is that the state of the ISI-IDS channel is defined by a few recent input bits, but the channel state in the Markov-IDS model is random and the distribution of the channel state is determined by the previous state. In ISI-IDS model, by using filler bits at the input, the total channel can be translated to multiple independent uses of the same channel. The Markov-IDS model, however, can only be translated to a product of similar (but not the same) independent channels. This made the analysis of the variance of mutual information density necessary.
After establishing the information stability of the Markov-IDS channels, we also provide some upper bounds on the channel capacity for a particular Markov deletion channel model. To do this, we employ the basic methodology developed in [23] by providing the transmitter and/or the receiver with suitable side information. Our numerical evaluations of the bounds indicate that having memory in the deletion process is beneficial in the sense that it increases the channel capacity for a given deletion probability (more precisely, upper bounds on the channel capacity, which are obtained in the same manner for both scenarios). In other words, we can quantify the expected rate increase for synchronization error channels with memory compared to those with i.i.d. synchronization errors.
The paper is organized as follows. In Section II, notations and information-theoretic definitions are introduced. In Section III, the Markov-IDS channel sequence is defined, and the capacity theorems are presented. In Section IV, the proofs of the capacity theorems are given. In Section V, the numerical upper bounds on the capacity for the Markov-IDS channel sequence are compared to those of the binary deletion channel. In Section VI, conclusions are drawn. In Appendix A, some helpful propositions about properties of functions of channels are drawn, which are used both in Dobrushin’s proofs and in the proofs in Section IV. In Appendix B, the variance of mutual information is analyzed via perturbation of the optimal input distribution.
II Preliminaries
A mathematical model of a physical channel is defined by conditional distributions $W(y|x)$ for all $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, which means that if a symbol $x\in{\mathcal{X}}$ is input, output $y\in{\mathcal{Y}}$ is observed with probability [^1] This is valid only for a countable ${\mathcal{Y}}$; for uncountable case, there should be output PDFs instead; in this paper uncountable ${\mathcal{Y}}$ are not considered. $W(y|x)$. The fact that $W$ is a channel with the input set ${\mathcal{X}}$, and the output set ${\mathcal{Y}}$ will be denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$. Shannon’s theorem [24] considers a memoryless channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ that is used $n$ times independently. The equivalent channel that transmits $n$ symbols simultaneously is the channel $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$, where the conditional probabilities are given by $W^{n}(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, with the use of the vector notation $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$.
In this paper, our focus is on channels with synchronization errors, modeled as insertions, deletions, and channel errors. As an example, consider an insertion channel $V_{n}$ with $n$ input symbols $x\in{\mathcal{X}}^{n}$. In general, for each input symbol $x_{i}\in{\mathcal{X}}$, an output vector $y_{i}$ over ${\mathcal{X}}$ of length $\geq 1$ is produced. In contrast with the channel $W^{n}$, the resulting output of $V_{n}$ is a concatenation of output vectors, but not the vector $y_{1}^{n}$ of output vectors. For example, if input bits are $(1,0,0)$ and the corresponding outputs for each bit are $(1,0)$, $(0)$ and $(1,0)$, then the overall channel output would be $(1,0,0,1,0)$, not $((1,0),(0),(1,0))$.
In general, a communication setup, besides conditional probabilities of the original channel, also includes some underlying method of combining the outputs of multiple channel uses. In the case of simple memoryless channels, the combining method is just stacking the output symbols into an output vector; in the case of insertion (and/or deletion) channel, the combining method is concatenation. For this reason, in the non-trivial setup a general case of a channel sequence $(W_{n}:{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{n})_{n\in{\mathbb{N}}}$ is usually considered [25, 26], where sets ${\mathcal{X}}_{n}$, ${\mathcal{Y}}_{n}$ can vary for different $n$. For simplicity of exposition, we consider the case when the input set for $W_{n}$ is ${\mathcal{X}}_{n}={\mathcal{X}}^{n}$ and ${\mathcal{X}}$ is fixed for all $n$.
For a channel sequence, the mutual information capacity is defined as
$\lim_{n\to\infty}{\mathbf{I}}(W_{n})/n,$ (1)
where ${\mathbf{I}}(W_{n})$ denotes the maximum mutual information between the input and the output of a channel $W_{n}$ over all possible input distributions. The coding capacity (or operational capacity) is the (maximum) asymptotic rate of coding schemes that achieve arbitrary small error probability as $n\to\infty$. The capacity theorem is a statement that the mutual information capacity is equal to the coding capacity.
There are many capacity theorems that generalize Shannon’s theorem on non-trivial channel sequences. Dobrushin’s work [27] considers the most general case, where input and output alphabets can be continuous, and the metric of the decoder output’s closeness to the original signal can be arbitrary. It is shown that the so-called “information stability” of a channel sequence is sufficient for the capacity theorem to hold. Later, in [25], it was shown that information stability is also necessary for the capacity theorem. In [9], Dobrushin proves the capacity theorem for channels with i.i.d. synchronization errors where an output vector of finite expected length is produced for each input symbol, and all the output vectors are concatenated to form the resulting output. The generalization of the capacity theorem to the case of IDS channels with continuous input alphabet is studied in [28]. Capacity theorems are also proven for FSMC [29], for Gilbert-Elliott channels [30], for deletion channels concatenated with FSMC [31], for IDS channels with intersymbol interference [22].
II-A Notation
@@
Sets are written in calligraphic letters ${\mathcal{X}},{\mathcal{Y}},{\overline{\mathcal{Y}}}$; operators are written in bold letters ${\mathbf{C}},{\mathbf{E}},{\mathbf{I}}$. For $b,c\in{\mathbb{N}}$, we denote a vector $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$. Also, we denote infinite sequences as $a^{\infty}=(a_{1},a_{2},\dots)$. For a set ${\mathcal{A}}$, we denote a set of all finite sequences with elements from ${\mathcal{A}}$ by ${\overline{\mathcal{A}}}={\left\{{()}\right\}}\cup\bigcup_{i=1}^{\infty}{\mathcal{A}}^{i}.$ We denote the length of a vector $a=a_{1}^{n}\in{\overline{\mathcal{A}}}$ by ${\mathbf{N}}(a)=n$. The concatenation (gluing) operator $\mathbf{g}$ is defined as $\mathbf{g}(a_{1}^{n},b_{1}^{k})=(a_{1},a_{2},...,a_{n},b_{1},...,b_{m}),$ which is also generalized to any number of input vectors. Also, define the function $\mathbf{g}_{n}:{\overline{\mathcal{A}}}^{n}\to{\overline{\mathcal{A}}}$, which glues $n$ vectors ${\overline{a}}_{i}\in{\overline{\mathcal{A}}}$ as $\mathbf{g}_{n}({\overline{a}}_{1}^{n})=\mathbf{g}({\overline{a}}_{1},{\overline{a}}_{2},...,{\overline{a}}_{n}).$ The inverse image $\mathbf{g}^{-1}_{n}(a)$ is the set of all possible ways of cutting $a\in{\overline{\mathcal{A}}}$ into $n$ subvectors, i.e., $\mathbf{g}^{-1}_{n}(a)={\left\{{{\overline{a}}_{1}^{n}|\mathbf{g}_{n}({\overline{a}}_{1}^{n})=a}\right\}}$. For a set ${\mathcal{S}}$, we denote by ${\mathbb{D}}_{\mathcal{S}}$ the set of all distributions over ${\mathcal{S}}$. For $n\in{\mathbb{N}}$, denote by ${\mathbb{D}}_{n}$ the set of distributions over the set ${\left\{{1,2,\dots,n}\right\}}$ in a form of row vectors $\rho_{1}^{n}$, $\sum_{i=1}^{n}\rho_{i}=1$, $\rho_{i}\geq 0$.
@@
The probability distribution for any probability, expectation, or entropy is written as a subscript of $\Pr$, ${\mathbf{E}}$, or ${\mathbf{H}}$. The values that are not known are interpreted as random variables, and we always take summation/expectation over all their possible values. For example, if we have a conditional distribution $W(u,x|v,y)$, where $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, and $u$ and $v$ are some fixed values, then
$\Pr_{W(u,x|v,y)}{\left[{x>0}\right]}$ $=\sum_{\begin{subarray}{c}x\in{\mathcal{X}}:x>0\\ y\in{\mathcal{Y}}\end{subarray}}W(u,x|v,y),$ (2)
${\mathbf{E}}_{W(u,x|v,y)}{\left[{f(x,y)}\right]}$ $=\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(u,x|v,y)\cdot f(x,y).$ (3)
@@
${\mathbf{H}}{\left[{W(x,y)}\right]}$ $=-\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(x,y)\cdot\log_{2}W(x,y).$ (4)
We use $O$- and $o$- notation for sequences / functions:
$a_{n}\in o(b_{n})$ $\iff\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0,$ (5)
$a_{n}\in O(b_{n})$ $\iff\lim_{n\to\infty}\frac{a_{n}}{b_{n}}\in{\mathbb{R}}.$ (6)
Note that $o(f(n))\subseteq O(f(n))$.